OpenStudy (anonymous):

Subtract and Simplify 7-v/v-6 - 3v-8/6-v =

OpenStudy (dumbcow):

-(6-v) = v-6

OpenStudy (anonymous):

Your notation might not be exactly precise, I think the problem is\[(7-v)/(v-6)-(3v-8)/(6-v)=(7-v+3v-8)/(v-6)=(2v-1)/(v-6)\]